Activity 2.3 Application of mathematical concepts and procedures in algebra

Mathematical idea (concept or procedure) Example of application in algebra

    1. Learnersmustbeabletoswitchcomfortablybetween all operations involving whole numbers and apply these operations.

Two ideas come together here:

• command of number and number sense

• comfortably operating numbers

In the Intermediate Phase these numbers are positive; in the Senior Phase learners start operating on integers (including positive and negative whole numbers).

  Do the following:

Calculate and write down the sum of the following:

• 11 + 12 + 13 + 14 + 15 =

• 23+24+25+26+27=200

+ 300 + 400 + 500 + 600 =

What is the fifth multiple of 13? What is the fifth multiple of 25? What is the fifth multiple of 400?

What is the rule that you can abstract from what you found above?

Explain why this is the case.


1
Expert's answer
2021-05-25T18:35:56-0400

1.

• 11 + 12 + 13 + 14 + 15 =65


• 23+24+25+26+27=125


• 200+ 300 + 400 + 500 + 600 =2000


Fifth Multiple of "13=13\times 5=65"


Fifth multiple of "25=25\times 5=125"


Fifth multiple of "400=400\times 5=2000"



We can conclude that Fifth multiple of a number is equal to the sum of the numbers from second predecessor to second successor. This is the case because When we calculate the fifth multiple It is always divisible by 5. and also We added the numbers in terms of 5 so we get the required form.